# Radio Map-Based Spectrum Sharing for Joint Communication and Sensing

### v1

@@ Abstract
The sixth-generation (6G) network is expected to provide both communication and sensing (C&S) services. However, spectrum scarcity poses a major challenge to the harmonious coexistence of C&S systems. Without effective cooperation, the interference resulting from spectrum sharing impairs the performance of both systems. This paper addresses C&S interference within a distributed network. Different from traditional schemes that require pilot-based high-frequency interactions between C&S systems, we introduce a third party named the radio map to provide the large-scale channel state information (CSI). With large-scale CSI, we optimize the transmit power of C&S systems to maximize the signal-to-interference-plus-noise ratio (SINR) for the radar detection, while meeting the ergodic rate requirement of the interfered user. Given the non-convexity of both the objective and constraint, we employ the techniques of auxiliary-function-based scaling and fraction programming for simplification. Subsequently, we propose an iterative algorithm to solve this problem. Simulation results collaborate our idea that the extrinsic information, i.e., positions and surroundings, is effective to decouple C&S interference.
@@ I INTRODUCTION
The advent of the fifth-generation (5G) network has ushered in a myriad of applications, ranging from automatic driving and smart manufacturing to telemedicine. These applications not only demand high communication data rates but also require precise sensing. The vision of enabling joint communication and sensing in the sixth-generation (6G) network has become increasingly clear. Driven by the ever-lasting demands, both C&S systems are increasing their scale in the full swing. However, their progress faces a significant obstacle of spectrum scarcity. The interference caused by spectrum sharing impairs the functionalities of both C&S systems [1, 2]. To coexist C&S systems and further build an integrated C&S network, JCAS has become a heated topic in both academia and industry.
In the literature, investigations on JCAS can be classified into two categories [3]. The first is C&S coexistence. In this case, C&S signals are generated and processed by independent C&S hardware. The C&S signals operate in the same frequency band, leading to mutual interference. Related studies mainly focus on the interference management [4, 5]. The second is C&S integration, which aims to integrate C&S functions into a single platform [6, 7]. In C&S integration, corresponding schemes can be further classified into three categories. The first is radar-centric. Related studies take sensing as the primary function and use radars as the dual-function platform [8, 9, 10]. The second is communication-centric. Related studies mainly focus on the data transmission and take the communication transceiver as the dual-function platform [11, 12]. The third category is novel waveform design, which aims to provide flexible C&S services using a new dual-function platform. Studies in this category are actively exploring the fundamental limits of C&S performance in a single transmission. Many dual-function waveforms have been proposed for different scenarios.
@@
Compared with C&S integration, which requires a long-term effort, the severe spectrum scarcity makes C&S coexistence a more urgent task at present. In the literature, many studies have delved into the interference mitigation between wireless communication and radar sensing. In the early stages, related studies were mainly unilateral, making adaptations either on the radar side or the communication side [13][22]. On the radar side, Bica et al. used mutual-information-based criteria to optimize the radar waveform [13]. Babaei et al. introduced a null-space precoder that confines radar signals within the null space of interference channels [14]. Shi et al. considered the imperfections of the channel state information (CSI) and proposed a power minimization scheme, with the C&S requirement serving as the constraints [15]. Kang et al. optimized the radar beampattern to manage its radiating energy in spatial directions and spectral frequency bands [16]. Alaee-Kerahroodi et al. proposed a cognitive radar waveform and validated its effectiveness using a custom-built prototype [17]. On the communication side, E. H. G. Yousif et al. devised a linear constraint variance minimization beamforming scheme to control the interference leakage from the base station (BS) to radars [18]. Liu et al. investigated the coexistence between a multi-input and multi-output (MIMO) radar and a multi-user MIMO communication system. They proposed a robust communication precoding scheme to maximize the detection probability while ensuring the minimal signal-to-interference-plus-noise ratio (SINR) of downlink users [19]. Moreover, the authors further paid attention to practical limitations in C&S cooperation [20]. They investigated the channel estimation and radar working mode judgment under different prior knowledge levels of the radar waveform. Wang et al. and Shtaiwi et al. introduced the reconfigurable intelligent surface (RIS) to the C&S network. Both of them investigated the joint BS and RIS precoding scheme to exploit the degrees of freedom (DoFs) of RIS for controlling C&S interference [21, 22].
@@
Due to the limited DoFs, the unilateral schemes are sometimes insufficient to control interference to a required level. The joint C&S cooperation is accepted as a more effective means [23][30]. Wang et al. considered a single-antenna C&S network and proposed a joint power allocation scheme from both radar-centric and communication-centric perspectives [23]. Li et al. investigated the matrix-completion-based radars in the coexistence design. With additional DoFs of subsampling of this kind of radar, they established an optimization framework that maximized the radar SINR with the constraint of the communication rate [24]. Qian et al. jointly optimized the radar transmit waveform, receive filter, and code book of the communication system in a clutter environment. To ensure the estimation accuracy, the authors took the waveform similarity as a constraint in the optimization [25]. Given that the interference generated by pulse radars is not constant, Zheng et al. proposed a novel metric named the compound rate to measure the communication performance under varying interference [26]. He et al. proposed a robust coexistence scheme with the objective of minimizing C&S power. The probability that C&S SINRs are below a certain threshold was constrained within a small value [27]. In addition to optimizing the detection performance, Cheng et al. minimized the Cramér-Rao Bound of the angle estimation by optimizing the radar waveform and communication transmit vectors [28]. Grossi et al. maximized the energy efficiency of the communication system while ensuring the received SINR of radars [29]. Qian et al. jointly optimized the radar receiver and communication transmitter to maximize the mutual information between the probing signals and received echoes [30].
@@
Although the studies mentioned above offered valuable insights into mitigating interference between C&S systems, a gap persists between the theoretical investigations and practical deployment. In many cases, practical conditions cannot support the perfect CSI assumption made in the literature. First, obtaining full CSI requires pilot-based channel estimation. It is nearly impossible for existing C&S systems to alter their hardware and software for receiving the counterpart signals. In addition, acquiring full CSI is costly. C&S systems have to send pilots, estimate CSI, and exchange CSI data during each channel coherence time, resulting in heavy signaling overhead. Moreover, C&S systems become less robust using the pilot-based cooperation. Any inaccuracies in CSI degrade cooperation efficiency, impacting the provisions of C&S services. Given these challenges, the pilot-based C&S cooperation may not be a practical choice in real-world applications.
@@
    We propose a radio-map-based C&S cooperation framework to mitigate C&S interference within a distributed network. Unlike traditional channel estimation relying on pilots, the radio map utilizes extrinsic environmental information to estimate the large-scale CSI. Leveraging this large-scale CSI, we propose a joint power allocation scheme that takes the radar SINR as the objective and the ergodic user rate as the constraint.
@@
    We employ a learning-based method to construct the radio map and utilize its predicted large-scale CSI to implement the proposed power allocation scheme. Simulation results validate the effectiveness of incorporating extrinsic information to coexist C&S systems in a loose cooperation manner.
@@ II SYSTEM MODEL
In Fig. 1, we illustrate the pilot-based C&S cooperation framework and the proposed radio-map-based C&S cooperation framework. As shown in Fig. 1 (a), the pilot-based C&S cooperation framework requires both signal-level and data-level C&S interactions. Within each channel coherence time, C&S systems exchange pilots, CSI data, and negotiate C&S requirements and allocation results. The two systems are tightly coupled by high-frequency and heavyweight interactions. In contrast, the radio-map-based framework only requires low-frequency and lightweight C&S interactions. Instead of transmitting pilots for CSI acquisition, C&S systems send original data to the radio map and get the large-scale CSI in return. With the radio-map-predicted large-scale CSI, C&S systems only need to negotiate their requirements and resource allocation results. Their negotiation frequency aligns with the varying frequency of the large-scale fading, which changes slowly. In this way, the burden of pilot-based online C&S interactions is alleviated through the offline interactions between the radio map and C&S systems, enabling C&S systems work in a loose cooperation manner.
Based on the proposed framework, we illustrate the specific system model of a radio-map-based distributed C&S network in Fig. 2. The communication system is composed of multiple distributed BSs and the number of BSs is denoted as $M_{c}$. These BSs are connected to a common processing center. They serve users in a user centric manner. To avoid inter-user interference, different users are allocated with the orthogonal frequency band. The sensing system is composed of multiple distributed radars and the number of radars is denoted as $M_{r}$. These radars operate in the same frequency band and use the orthogonally coded signals to tell from each other. Radars detect a common target. The detection results are sent to the fusion center to calculate the final judgment. In addition to C&S systems, a third party named the radio map is placed in the network. The radio map functions to exploit extrinsic environmental information to estimate the large-scale CSI. To facilitate the negotiation, optical fibers connect the communication processing center, radar fusion center and radio map. In this network, C&S systems share spectrum resources. We concentrate on one user whose allocated frequency band is overlapped with radars. The models and expressions can be extended to the case of multiple interfered users with few changes.
@@ II-A RADIO MAP MODEL
To specify the working regime of the radio-map-based C&S cooperation framework, we detail the inner modules and data flows of these three parties in Fig. 3. As depicted by the yellow blocks, the radio map includes two key modules: data management and radio map estimator. The data management module is responsible for data collection, storage and updates. It collects training data through self-measurement and from C&S systems. In the offline stage, the self-measurement approaches include the ray-tracing-based numerical calculation and on-site measurements. Simultaneously, the historical CSI of C&S systems can be sent to the radio map. In the online stage, the radio map can place dedicated sensors for real-time measurements [31]. Additionally, signals from C&S systems and radar detection results are also valuable real-time data. In each large-scale channel coherence time, the data management module processes the newly collected data and hands over them to the radio map estimator. Leveraging these data, the radio map estimator constructs or updates the estimator. From a mathematical perspective, the estimator is a mapping function, which maps the position data, i.e., the transmitting and receiving (T&R) positions, into the large-scale CSI [32]. In C&S systems, the data management modules collect T&R positions. They send T&R positions to the radio map and get the corresponding large-scale CSI back. In the depicted block diagram, the C&S resource management module resides within the radar fusion center. Consequently, the communication system needs to send its T&R positions to the radar fusion center, enabling the latter to know the large-scale CSI of both systems. With the large-scale CSI, the C&S resource management module runs the built-in algorithm, calculating the resource allocation results. These results are communicated to respective resource management modules and being implemented. It is worth noting that the C&S resource management module can alternatively be positioned within the communication processing center, integrated into the radio map, or function as an independent party to provide interference mitigation services. The specific configuration can be flexibly adapted to practical situations.
@@
Figure 3: Illustration of the working regime of the proposed radio-map-based C&S cooperation framework. The modules of the communication system, sensing system, and the radio map are depicted by the blue, green, and yellow blocks.
@@
Figure 4: Illustration of the lagre-scale CSI maps generated by Wireless Insite, the radio map and the curve fitting method. (a) Simulation Layout, (b) Real large-scale CSI calculated by Wireless Insite, (c) Radio-map-predicted large-scale CSI, and (d) Curve-fitting-method-predicted large-scale CSI. The large-scale CSI maps include one million data points each. These data are obtained by dividing the layout into one million grid points and calculating the channel gain between each grid point and the BS, #$\#1$, using the Wireless Insite, radio map and curve fitting method.
@@
To demonstrate the feasibility of the proposed framework, we present a case study of the radio map. A ray tracing software named Wireless Insite was introduced to calculate the large-scale CSI, which is regard as the ground truth data. In Fig. 4 (a), we illustrate the layout set in the software. It is a $5000$ m × $\times\ 5000$ m rectangular area, covering the range of $[-1500\ \text{m},3500\ \text{m}]$. There are randomly distributed buildings and lawns. In this simulation, we equally divided the layout into $4\times 10^{4}$ grids and placed a virtual receiver at each grid point. The software was used to calculate the channel gain between the transmitters (BSs/radars) and virtual receivers. We thereby obtained $4\times 10^{4}$ data points for each BS/radar. Using the $4\times 10^{4}$ data points and the corresponding positions of the virtual receiver, we trained a multi-layer perceptron network for each BS/radar. The network has five hidden layers of $[32,64,128,64,32]$ dimensions. It takes the receiving position as the input and outputs the channel gain. In Fig. 4 (c), we present the large-scale CSI map estimated by the radio map estimator of BS, #$\#1$. For comparison, the large-scale CSI map generated by Wireless Insite and estimated by the curve fitting method are also presented. The curve fitting method is based on a generic channel model:
@@
where $L$ represents the path loss, $d$ is the T&R distance, and $f_{c}$ is the carrier frequency. We applied the least squares method to calculate $\alpha$ and $\beta$ using the same training data as the radio map.
Through comparisons, we can see that the radio map provides a more accurate approximation to the real large-scale CSI compared to the curve fitting method. This is because the radio map is environmental-aware. By incorporating T&R positions, the radio map accounts for the losses caused by specific geographical factors of this link. It thus delivers more accurate large-scale CSI than generic empirical models. In addition, the large-scale CSI is quasi-static and slow-varying. For slow-moving scenarios, it is empirically sufficient for the radio map to collect data at the meter-level in the spatial domain and update the estimator at the second-level in the time domain [33]. The C&S interactions are also at the second level. Of course, the specific precision of the radio map and C&S interaction frequency depend on practical situations. Striking a balance between overhead and accuracy is crucial for ensuring the practical viability and effectiveness of the radio-map-based framework.

### v2

@@ Abstract
The sixth-generation (6G) network is expected to provide both communication and sensing (C&S) services. However, spectrum scarcity poses a major challenge to the harmonious coexistence of C&S systems. Without effective cooperation, the interference resulting from spectrum sharing impairs the performance of both systems. This paper addresses C&S interference within a distributed network. Different from traditional schemes that require pilot-based high-frequency interactions between C&S systems, we introduce a third party named the radio map to provide the large-scale channel state information (CSI). With large-scale CSI, we optimize the transmit power of C&S systems to maximize the signal-to-interference-plus-noise ratio (SINR) for the radar detection, while meeting the ergodic rate requirement of the interfered user. Given the non-convexity of both the objective and constraint, we employ the techniques of auxiliary-function-based scaling and fractional programming for simplification. Subsequently, we propose an iterative algorithm to solve this problem. Simulation results corroborate our idea that the extrinsic information, i.e., positions and surroundings, is effective to decouple C&S interference.
@@ I INTRODUCTION
The advent of the fifth-generation (5G) network has ushered in a myriad of applications, ranging from automatic driving and smart manufacturing to telemedicine. These applications not only demand high communication data rates but also require precise sensing. The vision of enabling joint communication and sensing in the sixth-generation (6G) network has become increasingly clear. Driven by the ever-lasting demands, both communication and sensing (C&S) systems are increasing their scale in full swing. However, their progress faces a significant obstacle of spectrum scarcity. The interference caused by spectrum sharing impairs the functionalities of both C&S systems [1, 2]. To coexist C&S systems and further build an integrated C&S network, the joint communication and sensing (JCAS) has become a heated topic in both academia and industry.
In the literature, investigations on JCAS can be classified into two categories [3]. The first is C&S coexistence. In this case, C&S functions are implemented using separate hardware modules. Due to the spectrum scarcity, C&S systems share spectrum resources. Consequently, they need to negotiate spectrum usage, power allocation, and temporal coordination to minimize interference. Additionally, the similar hardware structure and signal processing procedures of wireless communication and radio sensing have spurred the C&S integration. In this case, C&S functions are implemented by a common hardware. This integrative approach offers potential advantages in terms of cost-efficiency, reduced power consumption, and compactness and simultaneously facilitates a more versatile usage of radio resources. In the subsequent sections, we review related studies on C&S integration, followed by discussions on C&S coexistence.
Regarding C&S integration, corresponding studies can be subdivided into three categories. The first is the radar-centric approach, that uses radar platforms to transmit information [4, 5, 6, 7]. In the early stage, Hassanien et al. conducted several representative works, which modulated the radar sidelobe to achieve communication capabilities [4][5]. However, such schemes have a drawback of low data rates. Furthermore, the index modulation was incorporated into radar-centric designs. This technique uses a set of candidates, such as waveforms [6] and transmit codes [7], and embeds the information into the permutation and combination of these elements. Due to the vast number of possible permutations and combinations, the system can transmit data at Mbps rates without substantial modification of the primary radar signal. The second is the communication-centric approach. Related studies mainly investigated signal processing schemes to extract information from communication signals. For example, Kumari et al. applied the preamble of IEEE 802.11ad for the target detection and parameter estimation, demonstrating that centimeter-level accuracy is achievable thanks to the perfect auto-correlation property of the preamble [8]. Pucci et al. explored the sensing capabilities of the orthogonal frequency division multiplexing (OFDM) waveforms. The authors derived the expressions of the detection probability and the root mean squared error of the position and velocity estimation [9]. The third category is the novel waveform design, which aims to provide flexible C&S services using a new dual-function waveform. For example, Sturm et al. discussed the basic requirements of the dual-function waveform and the processing techniques to extract information [10]. Liu et al. proposed a solution to balance C&S by minimizing a weighted objective of inter-user interference and deviations from constant-modulus waveforms. The authors also developed a low-complexity algorithm to facilitate practical implementations [11]. Xiao et al. proposed a full-duplex JCAS platform which leverages the idle window of probing pulses to transmit communication bits. The proposed scheme achieved a higher data rate than radar-centric schemes and better sensing performance than communication-centric schemes [12].
@@
Compared with C&S integration, which requires a long-term effort, the severe spectrum scarcity makes C&S coexistence a more urgent task at present. In this paper, we thus concentrate on the interference mitigation between C&S systems. Related works are reviewed as follows. In the early stage, related studies were mainly unilateral which optimized either the radar side or the communication side [13][25]. On the radar side, Bica et al. used mutual-information-based criteria to optimize the radar waveform [13]. Babaei et al. introduced a null-space precoder that confines radar signals within the null space of interference channels [14]. Shi et al. considered the inaccuracy of channel state information (CSI) and proposed a power minimization scheme, with the C&S requirement serving as the constraints [15]. Kang et al. optimized the radar beampattern to manage its radiating energy in spatial directions and spectral frequency bands [16]. Alaee-Kerahroodi et al. proposed a cognitive radar waveform and validated its effectiveness using a custom-built prototype [17]. On the communication side, E. H. G. Yousif et al. devised a linear constraint variance minimization beamforming scheme to control the interference leakage from the base station (BS) to radars [18]. Raymond et al. utilized a spatio-temporal analytical method to quantify the instantaneous and average interference suffered by radar from neighboring BSs. Building upon this analysis, they introduced an adaptive BS power control strategy that simultaneously safeguards primary radar operations while ensuring effective cellular communication [19]. Singh et al. explored spectrum sharing between a full-duplex multi-input and multi-output (MIMO) BS and a MIMO radar. The authors proposed a power allocation scheme which takes the detection probability as the objective, subject to the signal-to-interference-plus-noise ratio (SINR) requirements for both uplink and downlink communications [20]. Labib et al. examined the coexistence between the radar and LTE system. The authors proposed a stochastic optimization scheme to maximize the BS transmit power within the bound of acceptable interference at the radar side [21]. Liu et al. investigated the coexistence between a MIMO radar and a multi-user MIMO communication system. They proposed a robust communication precoding scheme to maximize the detection probability while ensuring the minimal SINR of downlink users [22]. Moreover, the authors further paid attention to practical limitations in C&S cooperation [23]. They investigated the channel estimation and radar working mode judgment under different levels of prior knowledge of the radar waveform. Recently, Wang et al. and Shtaiwi et al. introduced the reconfigurable intelligent surface (RIS) to the C&S network. Both of them investigated the joint BS and RIS precoding scheme to exploit the degrees of freedom (DoFs) of RIS for controlling C&S interference [24, 25].
@@
Due to the limited DoFs, the unilateral schemes sometimes are insufficient to control interference effectively. The joint C&S cooperation is accepted as a more effective means [26][38]. Wang et al. considered a single-antenna C&S network and proposed a joint power allocation scheme from both radar-centric and communication-centric perspectives [26]. Furthermore, the authors extended the power allocation scheme to multi-carrier C&S systems with additional consideration of the clutters [27]. Martone et al. proposed a multi-objective optimization scheme, which adjusts power, frequency, and bandwidth parameters to maximize both C&S criteria [28]. Considering the competing relationships between C&S systems, Mishra et al. modeled C&S interactions as a two-person zero-sum game and proposed a power allocation strategy to find the equilibrium between C&S systems [29]. Tian et al. applied mutual information as the metric of radar sensing and proposed a power allocation scheme that simultaneously maximizes mutual information of radar sensing and communication data rate [30]. Li et al. investigated the matrix-completion-based radars in the coexistence design. With additional DoFs of subsampling of this kind of radar, they established an optimization framework that maximized the radar SINR with the constraint of the communication rate [31]. Rihan et al. investigated the spectrum sharing between MIMO radar and MIMO communication systems. The authors proposed a two-tier alternating optimization scheme based on the interference alignment approach, which optimizes transmit precoders and receive spatial filters to maximize both the radar SINR and communication sum rate [32]. Qian et al. jointly optimized the radar transmit waveform, receive filter, and code book of the communication system in a clutter environment. To ensure the estimation accuracy, the authors took the waveform similarity as a constraint in the optimization [33]. Given that the interference generated by pulse radars is not constant, Zheng et al. proposed a novel metric named the compound rate to measure the communication performance under varying interference [34]. He et al. proposed a robust coexistence scheme with the objective of minimizing C&S power. The probability that C&S SINRs are below a certain threshold was constrained within a small value [35]. In addition to optimizing the detection performance, Cheng et al. minimized the Cramér-Rao Bound of the angle estimation by optimizing the radar waveform and communication transmit vectors [36]. Grossi et al. maximized the energy efficiency of the communication system while ensuring the received SINR of radars [37]. Qian et al. jointly optimized the radar receiver and communication transmitter to maximize the mutual information between the probing signals and received echoes [38].
@@
Although the studies mentioned above offered valuable insights into mitigating interference between C&S systems, a gap persists between the theoretical investigations and practical deployment. In many cases, practical conditions cannot support the perfect CSI assumption used in the literature. First, obtaining full CSI requires pilot-based channel estimation. It is nearly impossible for existing C&S systems to alter their hardware and software for receiving the counterpart signals. In addition, acquiring full CSI is costly. C&S systems have to send pilots, estimate CSI, and exchange CSI data during each channel coherence time, resulting in heavy signaling overhead. Moreover, C&S systems are not robust using the pilot-based cooperation. The inaccuracy of CSI would degrade cooperation efficiency, impacting the provisions of C&S services. Given these challenges, the pilot-based C&S cooperation maybe not a practical choice in real-world applications.
@@
    We propose a radio-map-based C&S cooperation framework to mitigate C&S interference in a distributed network. Unlike traditional channel estimation relying on pilots, the radio map utilizes extrinsic environmental information to estimate the large-scale CSI. Leveraging the large-scale CSI, we propose a joint power allocation scheme that takes the radar SINR as the objective and the ergodic user rate as the constraint.
@@
    We employ a learning-based method to construct the radio map and utilize its predicted large-scale CSI to implement the proposed power allocation scheme. Simulation results validate the effectiveness of incorporating extrinsic information to coexist C&S systems in a loosely cooperated manner.
In addition, this work represents an extension of our previous work [39]. Unlike [39], which assumed the radio map is ready-made, this work unites the map construction and usage to provide a complete radio-map-based framework. We enhance the mathematical work with detailed proofs and conduct unified simulations by incorporating the radio map and proposed power allocation scheme. More in-depth discussions and insights are also provided throughout the paper. Therefore, this work is a necessary extension and enhancement of our previous work.
@@ II SYSTEM MODEL
In Fig. 1, we illustrate the pilot-based C&S cooperation framework and the proposed radio-map-based C&S cooperation framework. As shown in Fig. 1 (a), the pilot-based C&S cooperation framework requires both signal-level and data-level interactions between C&S systems. Within each channel coherence time, C&S systems exchange pilots, CSI data, C&S requirements, and allocation results. The two systems are tightly coupled by high-frequency and heavyweight interactions. In contrast, the radio-map-based framework only requires low-frequency and lightweight C&S interactions. Instead of using pilots to estimate channel conditions, C&S systems send original data to the radio map and get the large-scale CSI back. With the radio-map-predicted large-scale CSI, C&S systems only need to negotiate their requirements and resource allocation results. Their negotiation frequency is the same as the varying frequency of the large-scale fading, which changes slowly. C&S systems thus work in a loosely cooperated manner.
Based on the proposed framework, we illustrate our radio-map-based distributed C&S network in Fig. 2. The communication system is composed of $M_{c}$ distributed BSs. These BSs are connected to a common processing center. They serve users in a user-centric manner. To avoid inter-user interference, different users are allocated with orthogonal frequency bands. The sensing system is composed of $M_{r}$ distributed radars. These radars operate in the same frequency band and use the orthogonal signals to tell from each other. Radars detect a common target. The detection results are sent to the fusion center to calculate the final decision. In addition to C&S systems, a third party named radio map is placed in the network to estimate the large-scale CSI. To facilitate the negotiation, optical fibers connect the communication processing center, radar fusion center, and radio map. In this network, C&S systems share spectrum resources. We concentrate on one user whose allocated frequency band is overlapped with radars. The models and expressions can be extended to the case of multiple interfered users with few changes.
@@ II-A RADIO MAP MODEL
In Fig. 3, we illustrate the working regime of the radio-map-based C&S cooperation framework. As depicted by the yellow blocks, the radio map includes two key modules: data management and radio map estimator. The practical implementation of the radio map involves three steps: data collection and map construction, integration with C&S systems, and map update. Initially, the data management module extensively collects data. These data are collected by the approaches such as numerical calculations using the ray-tracing method and on-site measurements [40]. The original data generated by C&S systems in their daily operations can also be sent to the radio map. The data are preprocessed by the data management module and handed over to the radio map estimator, which is responsible for constructing the radio map. In the literature, the map construction methods can be divided into two categories: model-based and data-based. Model-based methods are based on a prior model and use collected data to estimate model parameters. Conventional channel models are a kind of model-based radio maps. These channel models are broadly categorized according to the transmission environment, e.g., urban, sub-urban, rural, and sub-rural [41], and the model parameters are estimated using the data collected from corresponding environments. However, due to the rough division, radio maps generated from model-based methods cannot accurately describe specific environments. Data-based methods, on the other hand, use data exclusively to construct the radio map. Related data-learning methods include kernel-based learning, kriging, dictionary learning, and deep learning [42]. These radio maps are environmental-aware and capable of providing accurate radio information for the areas where the data are collected. However, data-based methods suffer from the high cost of data management and high computational complexity [43]. Therefore, some studies begin to explore combining these two methods in the construction [44][45]  [^1] In this paper, the radio map exclusively refers to the map that is constructed for a certain area and is environmental-aware. . Afterwards, the second step is to integrate the radio map into C&S systems. The radio map provider needs to negotiate with C&S systems on the specifics such as the required original data, the provided information and the frequency of this provision. The third step involves the map update. To reflect the dynamic environment, the data management module persistently gathers and processes data. The newly collected data are periodically relayed to the radio map estimator, which updates the radio map to capture the dynamics of the real world.
@@
In our framework, the radio map works to provide the large-scale CSI. It maps position data, i.e., the transmitting and receiving (T&R) positions, into large-scale CSI [46]. This position information is collected by the data management modules of C&S systems. They send T&R positions to the radio map and get the corresponding large-scale CSI back. With the large-scale CSI, the C&S resource management module runs the built-in algorithm, calculating the resource allocation results. Then, these results are sent to resource management modules of C&S systems and being implemented. In the presented figure, the C&S resource management module resides in the radar side. Consequently, the communication system sends its T&R positions to the radar fusion center, enabling the C&S resource management module to know the large-scale CSI of both systems. The C&S resource management module can alternatively be positioned within the communication processing center, integrated into the radio map, or function as an independent party to provide interference mitigation solutions. The specific configuration can be flexibly adapted to practical situations.
@@
Figure 3: Illustration of the working regime of the proposed radio-map-based C&S cooperation framework. The modules of the communication system, sensing system, and the radio map are depicted by the blue, green, and yellow blocks, respectively.
@@
Figure 4: Illustration of the large-scale CSI maps generated by Wireless Insite, the radio map and the curve fitting method. (a) Simulation layout, (b) MLP structure (c) Real large-scale CSI calculated by Wireless Insite, (d) Radio-map-predicted large-scale CSI, and (e) Curve-fitting-method-predicted large-scale CSI. The MLP consists of a two-dimensional input layer, one-dimensional output layer, and five hidden layers with [32,64,128,64,32] neurons. The large-scale CSI maps include one million data points each. These data are obtained by dividing the layout into one million grid points and calculating the channel gain between each grid point and the BS, #$\#1$, using the Wireless Insite, radio map and curve fitting method.
@@
To demonstrate the feasibility of the proposed framework, we present a case study of the radio map. A ray tracing software named Wireless Insite was introduced to calculate the large-scale CSI, which is regarded as the ground truth data. In Fig. 4 (a), we illustrate the layout set in the software. It is a $5000$ m × $\times\ 5000$ m rectangular area, covering the range of $[-1500\ \text{m},3500\ \text{m}]$. There are randomly distributed buildings and lawns. In this simulation, we equally divided the layout into $4\times 10^{4}$ grids and placed a virtual receiver at each grid point. The software was used to calculate the channel gain between the transmitters (BSs and radars) and virtual receivers. We thereby obtained $4\times 10^{4}$ data points for each BS and radar. Using the $4\times 10^{4}$ data points and the corresponding positions of the virtual receiver, we trained a multi-layer perceptron (MLP) for each BS and radar. The structure of the MLP is presented in Fig. 4 (b). During data preprocessing, we applied the max-min normalization to scale the data within the range of $[-1,1]$. Our MLP takes the normalized receiving position, i.e, $[\bar{x},\bar{y}]$, as the input and takes the normalized channel gain, i.e., $\bar{l}$, as the output. The MLP has five hidden layers with $[32,64,128,64,32]$ neurons and between these layers the hyperbolic tangent function is used as the activation function. For training, we employed the Adaptive Moment Estimation (ADAM) optimizer and used the default parameters as defined in TensorFlow: learning rate of $0.001$, $\beta=(0.9,0.999)$, and $\epsilon=1e-8$ [47]. The network was trained over $6000$ epochs with a batch size of $128$, and the mean square error (MSE) was the loss function. Additionally, the learning rate was halved every $2000$ epochs to refine the learning process. In Fig. 4 (c), we present the large-scale CSI map estimated by the radio map estimator of BS, #$\#1$. For comparison, large-scale CSI maps generated by Wireless Insite and estimated by the curve fitting method are also presented. The curve fitting method is based on a generic channel model [48]:
@@
where $L$ represents the path loss, $d$ is the T&R distance, and $f_{c}$ is the carrier frequency. We applied the least squares method to calculate $\alpha$ and $\beta$. The used data are the same as those used for constructing the radio map.
Through comparisons, we can see that the radio map provides a more accurate approximation to the real large-scale CSI compared to the curve fitting method. This is because the radio map is environment-aware. By incorporating T&R positions, the radio map considers the losses caused by specific geographical factors along this link. It thus delivers more accurate large-scale CSI than generic empirical models. In addition, we discuss the applicability of the proposed framework for different moving-speed targets. As shown in Fig. 3, the update frequency of the large-scale CSI determines the interaction frequency among the radio map and C&S systems. In the scenario of slow-moving targets, the large-scale CSI is slow-varying. As a result, the radio map and C&S systems only need low-frequency interactions. While in the contexts with high-speed targets, C&S systems need high-frequency interactions to update the target position and large-scale CSI. This lets C&S systems to be tightly coupled. In addition, to adequately cover the area of target activities, the radio map needs to be large, which greatly increases its construction and maintenance costs. The benefits of our radio-map-based framework over traditional pilot-based framework become less pronounced. Therefore, this paper primarily focuses on the applicability of our radio-map-based framework in scenarios with slow-moving targets. Potential extension to fast-moving targets necessitates the development of cost-effective algorithms for large-scale radio map construction and update.
Moreover, the radio map, recording the information of the electromagnetic environment, its usage is more than C&S coordination. With enhanced environmental awareness, the radio map contributes to the efficiency of networks, facilitating a smarter utilization of the spectrum resources. For example, radio maps help in dynamic spectrum management by identifying spectrum usage patterns and detecting interference sources. They also assist in urban sensing by providing electromagnetic profile data, which help determine the location and size of buildings. Additionally, radio maps play an active role in localization, enabling accurate positioning by comparing real-time signal measurements with pre-constructed radio maps. In security applications, radio maps assist in detecting unauthorized devices and monitoring signal activities within a specified area. Radio maps can also be applied in environmental monitoring, such as assessing air conditions by providing signal propagation attenuation data. By incorporating the radio map with other advanced technologies, e.g., machine learning and Internet of things (IoT) [49], the sensing accuracy, communication efficiency, and overall intelligence will be greatly enhanced. However, realizing the full potential of radio maps necessitates dedicated research efforts in several key areas. Fast and efficient map construction and update are critical to ensure the radio map remain effective for the dynamic electromagnetic environment. Seamless integration with C&S systems requires the development of interoperable standards and protocols. Moreover, striking a balance between the costs and benefits of the radio map is paramount to ensure its widespread deployment. We call for more research efforts to explore these promising avenues further.